# NILE [Uncaptioned image] : Internal Consistency Alignment in Large Language Models

### v1

@@ 4 Experiments
For the main experiment, we use open source models like Mistral-7b-v0.3 Jiang et al. (2023) and Llama-3.1-8b Dubey et al. (2024) on two public datasets Alpaca Taori et al. and OpenOrca Mukherjee et al. (2023) to examine NILE framework’s robustness extensively. In addition, we conduct an ablation study to evaluate the efficacy of our design choices in the pipeline. More experiment details and case studies can be found in Appendix A.1.
@@ 4.1 IFT Datasets
The Alpaca dataset contains 52,000 instruction-following data generated using the techniques in the Self-Instruct Wang et al. (2023b). It starts with a limited (e.g., $175$ in our study) seed set of manually written tasks that are used to guide the overall generation. Then language models are utilized and prompted to augment these instructions and create corresponding instruction-answer instances. In our experiments, we use all the samples in a newer version of Alpaca [^1] https://huggingface.co/datasets/vicgalle/alpaca-gpt4 dataset, which includes instruction-following instances generated using GPT-4 Peng et al. (2023).
@@ 4.2 Evaluation
Alpaca-Eval V2
Alpaca-Eval V2444https://github.com/tatsu-lab/alpaca_eval is an automatic evaluation system for instruction-following language models Dubois et al. (2024). It builds upon the original AlpacaEval system, which benchmarked against OpenAI’s Davinci-003. AlpacaEval V2 instead uses GPT-4-Turbo, signaling the new state-of-the-art model since the original system’s creation.
@@ 4.3 Implementation details
We finetune Mistra-7b-v0.3 and Meta-Llama-3.1-8b models for our experiments. For selecting retriever $\mathcal{R}$ in IKE, we find that BM25 is more effective than a strong neural retriever such as contriver Lei et al. (2023) in retrieving higher-quality demonstrations, which is evaluated and validated in Appendix A.1.2. To maintain a better state of internal consistency, we set $\beta$ in Eq. 3 to the 1-st percentile of the dataset for Alpaca, and to the 2-nd percentile for OpenOrca to rule out a small amount of low ICI samples.
@@ 4.5 Results on Orca Dataset
Table 4 shows the performance of our NILE framework and all baselines on model Mistral-7b-v0.3 and Llame-3.1-8b in OpenOrca dataset. As we can see, Orca dataset brings unbalanced improvements on different LLMs, with LLama-3 having less improvements on Arena-Hard and Alpaca-Eval V2 LCWR and more on MTBench and BBH than Mistral, which reflect different underlying characteristics and potentially difference internal knowledge in these two models.
@@ 4.6 Results on Alpaca
Compared with Orca dataset, LLMs finetuned with Alpaca dataset is generally weaker than ones with Orca, which highlights the sheer quality differences between the two datasets. Desipte these differences, Alpaca + NILE still bring signficant improvements over Alpaca vanilla in all metrics, coming close to or even surpassing Orca vanilla in most of the benchmarks except BHH. It achieves a performance uplift by 3.7 and 4.1 points on Alpaca-Eval V2 LCWR and BBH in Mistral. Moreover, Alpaca + NILE raises Alpaca-Eval V2 LCWR and Arena-Hard by 3.1 and 2.7 in Llama-3.
@@ 4.7 Ablation Study
4.7.1 Effects of Different Internal Knowledge Sources
@@
Alpaca + KSR (Mistral) 4.00 9.14 / 7.29 6.64 57.67
Alpaca + KSR (Llama) 4.80 10.75 / 9.38 6.67 60.73
Orca + KSR (Mistral) 5.10 12.50 / 10.25 5.93 22.32
Orca + KSR (Llama) 5.20 13.67 / 11.21 7.51 64.03
Table 5: Effects of KSR in Llama-3 finetuning with internal knowledge from different LLMs. The highest values are bolded.
4.7.2 Effects of IKE Fewshot Number
Table 6 examines how different few-shot numbers of demonstration learning in IKE affect LLM performance. Here we evaluate three variants: 1) w. FD, which extracts LLM’s internal knowledge with a fixed 2-shot demonstrations described in A.2.2; 2) w. FS 1 IKE, which retrieves the top 1 most similar samples with BM25 as demonstrations; and 3) w. FS 2 IKE, which retrieves the top 2 most similar samples with BM25 as demonstrations; Though w. FS 2 IKE leads to degradation in some benchmarks, such as Arena-Hard for Alpaca and Alpaca-Eval for Orca, it still achieves overall improvements with BBH for Alpaca increasing by 0.7 and Arena-Hard for Orca increasing by 0.3. The results show the IKE phase is necessary for unaligned LLMs to more effectively extract internal knowledge, while fixed prompting reaches subpar performance.
Alpaca + KSR w. FD 4.80 10.75 / 9.38 6.67 60.73
Alpaca + KSR w. FS 1 IKE 4.50 11.20 / 9.75 6.72 59.25
Alpaca + KSR w. FS 2 IKE 4.50 10.82 / 10.56 6.76 61.40
Orca + KSR w. FD 5.20 13.67 / 11.21 7.51 64.03
Orca + KSR w. FS 1 IKE 4.90 12.46 / 10.99 7.40 63.89
Orca + KSR w. FS 2 IKE 5.50 13.00 / 11.50 7.43 64.29
Table 6: Effects of IKE with different fewshot numbers (FS) in Llama-3. The highest values are bolded, and the second highest is underlined.
4.7.3 Effects of KAR
@@
Alpaca + NILE wo. ICF 4.50 10.82 / 10.56 6.76 61.40
Alpaca + NILE w. ICF (low) 4.80 10.69 / 10.43 6.90 61.40
Alpaca + NILE w. ICF (high) 4.50 9.92 / 9.70 6.79 61.71
Orca + NILE wo. ICF 5.50 13.00 / 11.50 7.43 64.29
Orca + NILE w. ICF (low) 6.00 13.70 / 12.11 7.48 64.05
Orca + NILE w. ICF (high) 4.80 13.19 / 11.49 7.30 63.95
Table 7: Effects of ICF using different $\beta$ parameters in Llama-3. The highest values are bolded, and the second highest is underlined.
To validate KAR’s effectiveness in enhancing internal consistency between world knowledge from instructions and the model’s internal knowledge, we conducted experiments measuring the similarity between extracted internal knowledge and baseline knowledge across different models and datasets. In Llama-3 and Mistral, we used the instructions from the Alpaca and Orca as prompts to evaluate the models’ internal knowledge. Then, we obtained the models’ vanilla output for these instructions, the output generated using KSR, and the output using SR. We randomly sampled 10,000 instructions to calculate the sentence similarity between three generated outputs and the internal knowledge. As demonstrated in Figure 3, the similarity score distribution of the outputs generated by KSR is significantly closer to 1 compared to the other two methods, with Chi-squared test p-values lower than 0.01. More experiment results are exhibited in Appendix A.2.3. These results provide strong evidence supporting the effectiveness of the KSR approach in increasing the internal consistency from instructions by integrating relevant world and internal knowledge.
4.7.4 Effects of ICF
Table 7 looks into the effect of ICF. $\beta$ is set to 1-st percentile in Alpaca + NILE (low) and to 2-nd percentile in Orca + NILE (low). For both Alpaca + NILE (high) and Orca + NILE (high), we set $\beta$ to 10-th percentile. The results empirically prove that striking a balance between consistent and inconsistent knowledge in the IFT dataset is necessary for NILE to achieve ideal performance. We find the general advantage of ALPACA + NILE (low) over ALPACA + NILE (high) and discarding ICF (ALPACA + NILE wo. ICF), indicating that a surplus of overly consistent or inconsistent samples in IFT datasets both hurt LLM’s performance, and it is crucial to find the middle ground in these samples. This experiment further verifies our design choices of the ICF phase in NILE.
@@ 5 Conclusion
We present NILE, an innovative framework designed to enhance the quality of IFT datasets by aligning them with LLMs’ internal knowledge. Our extensive experiments demonstrate substantial improvements across various benchmarks, highlighting the crucial role of maintaining consistency between models’ internal knowledge and external knowledge in datasets. Each component of the NILE framework has been validated, reinforcing its importance in achieving better alignment. NILE offers promising directions for boosting the capabilities of LLMs and unlocking their full potential.
@@
While NILE can already obtain satisfactory performance, future works should expand NILE’s training by utilizing the complete OpenOrca dataset rather than the current 50,000-sample subset (5% of the dataset), due to limited time and computational resources. Additionally, future research should examine NILE’s capability for iterative instruction refinement, as the current implementation uses only a single revision pass. These expansions could further enhance NILE’s instruction-following capabilities.
@@
We conduct this study strictly under the guidance of community ethical principles. The utilized IFT datasets are reported to be safe and free of content that may contain discrimination, personally identifiable information, or any other undesirable behaviors. We meticulously curate our instructions to the LLMs to ensure that the tasks are limited to knowledge generation and knowledge-relevant revisions, thereby avoiding content that may pose ethical concerns.
@@ A.1 Implementation Details of NILE
For all experiments in this work, we use the Python 3.10.14 environment and vLLM 0.5.5 library [^7] https://github.com/vllm-project/vllm for LLM local inference of both Mistral and Llama-3. For vLLM inference hypermeters, we set the random seed to 42, max_tokens to 1024, temperature to 0.7, top_k to 50, top_p to 0.7, and repetition_penalty to 1. We run all experiments on a server with an Intel Xeon Silver 4309Y CPU and 8 Nvidia RTX A6000 GPU having 48GB GDDR6 VRAM, and we utilize official checkpoints Mistral-7B-v0.3888https://huggingface.co/mistralai/Mistral-7B-v0.3 for Mistral and Llama-3.1-8B [^9] https://huggingface.co/meta-llama/Llama-3.1-8B for Llama-3. For LLM instruction fine-tuning in this work, we choose llama-recipes [^10] https://github.com/meta-llama/llama-recipes for Llama-3 and alignment-handbook [^11] https://github.com/huggingface/alignment-handbook for Mistral. For Llama-3 fine-tuning, we set context_length to 2048, gradient_accumulation_step to 32, learning rate to 2e-5, and training batch size to 4. As for Mistral fine-tuning, we set context_length to 2048, gradient_accumulation_step to 32, learning rate to 2e-5, training batch size to 4, lr_scheduler_type to "cosine", num_train_epochs to 3, and warmup_ratio to 0.1. Fine-tuning for both Llama-3 and Mistral is done within 5 hours using 8 A6000 GPU.
@@ A.1.1 Internal Knowledge Extraction (IKE)
For demonstration sample, we randomly sample $m=5,000$ instruction pairs $q^{d}_{i}=\{\mathrm{instruction}^{d}_{i},\mathrm{input}^{d}_{i}\}$ from Alpaca dataset [^12] https://huggingface.co/datasets/vicgalle/alpaca-gpt4, since instructions in it are simple and straightforward, which is suitable for LLM demonstration learning. We leverage gpt-4-turbo-2024-04-09 through Openai API as the teacher LLMs for generating demonstrations given $q^{d}_{i}$ shown in Table 1. For GPT-4 endpoints, we use openai 1.42.0 python library and set n to 1, temperature to 0.7, and max_tokens to 1,024. We stick to the regulations from the OpenAI company when accessing its API. For retriever $\mathcal{R}$ in IKE, we choose the BM25 and Contriver implementation [^13] https://github.com/castorini/pyserini from the Pyserini 0.38.0 library.

### v2

@@ 4 Experiments
For the main experiment, we use open source models like Mistral-7b-v0.3 Jiang et al. (2023) and Meta-Llama-3.1-8b Dubey et al. (2024) on two public datasets Alpaca Taori et al. and OpenOrca Mukherjee et al. (2023) to examine NILE framework’s robustness extensively. In addition, we conduct an ablation study to evaluate the efficacy of our design choices in the pipeline. More experiment details, ablation study, inference overhead, and case studies can be found in  A.1.
@@ 4.1 IFT Datasets
The Alpaca dataset contains 52,000 instruction-following data generated using the techniques in the Self-Instruct Wang et al. (2023c). It starts with a limited (e.g., $175$ in our study) seed set of manually written tasks that are used to guide the overall generation. Then language models are utilized and prompted to augment these instructions and create corresponding instruction-answer instances. In our experiments, we use all the samples in a newer version of Alpaca [^2] https://huggingface.co/datasets/vicgalle/alpaca-gpt4 dataset, which includes instruction-following instances generated using GPT-4 Peng et al. (2023).
@@ 4.2 Evaluation
Alpaca-Eval V2 (A.-E. V2)
Alpaca-Eval V2555https://github.com/tatsu-lab/alpaca_eval is an automatic evaluation system for instruction-following language models Dubois et al. (2024). It builds upon the original AlpacaEval system, which benchmarked against OpenAI’s Davinci-003. AlpacaEval V2 instead uses GPT-4-Turbo, signaling the new state-of-the-art model since the original system’s creation.
@@ 4.3 Implementation details
For our experiments, we fine-tune the pretrained but unaligned models, Mistral-7b-v0.3 and Meta-Llama-3.1-8b. For selecting retriever $\mathcal{R}$ in IKE, we find that BM25 is more effective than a strong neural retriever such as contriver Lei et al. (2023) in retrieving higher-quality demonstrations, which is evaluated and validated in Appendix A.1.2. To maintain a better state of internal consistency, we set $\beta$ in Eq. 3 to the 1st percentile of the ICI distribution for Alpaca and to the 2nd percentile for OpenOrca to rule out a small amount of low ICI samples. Based on our manual random screening of 100 sample points respectively in Alpaca and OpenOrca datasets, we found the selected values in ICF to be a reasonable balance - lower thresholds retain too many misaligned knowledge samples that could directly impair performance, while higher thresholds discard too many consistent samples.
@@ 4.4 Baselines
To maintain consistency and a fair comparison with the Vanilla setting, the implementation of NILE and SR baseline rewrites only the responses $a^{o}$, leaving the rest of the dataset unchanged.
@@ 4.5 Results on Orca Dataset
Table 4 shows the performance of our NILE framework and all baselines on model Mistral-7b-v0.3 and Meta-Llama-3.1-8b in OpenOrca dataset. As we can see, Orca dataset brings unbalanced improvements on different LLMs, with LLama-3 having less improvements on Arena-Hard and Alpaca-Eval V2 LCWR and more on MTBench and BBH than Mistral, which reflects different underlying characteristics and potentially distinct internal knowledge in these two models.
@@ 4.6 Results on Alpaca
Compared with Orca dataset, LLMs finetuned with Alpaca dataset are generally weaker than ones with Orca, which highlights the sheer quality differences between the two datasets. Despite these differences, Alpaca + NILE still brings significant improvements over Alpaca vanilla in all metrics, coming close to or even surpassing Orca vanilla in most of the benchmarks except BBH. It achieves a performance uplift by 3.7 and 4.1 points on Alpaca-Eval V2 LCWR and BBH in Mistral. Moreover, Alpaca + NILE raises Alpaca-Eval V2 LCWR and Arena-Hard by 3.1 and 2.7 in Llama-3.
@@
4.7 Experiment Results on More LLMs
We conduct additional experiments of NILE on pretrained models of varying sizes, such as Meta-Llama-3.2-3B [^8] https://huggingface.co/meta-llama/Llama-3.2-3B, as well as on different model families, including Qwen2.5-7B [^9] https://huggingface.co/Qwen/Qwen2.5-7B and Qwen2.5-14B101010hhttps://huggingface.co/Qwen/Qwen2.5-14B, using the MTBench and Alpaca-Eval V2 benchmarks. The consistent and significant relative improvements (up to 85.7% on Alpaca-Eval V2 LCWR) observed in Table 5 demonstrate that NILE consistently delivers meaningful improvements across diverse LLM configurations.
Method MTBench $\uparrow$ Alpaca-Eval V2 $\uparrow$
Meta-Llama-3.2-3B
Alpaca vanilla 5.52 6.17 / 3.54
Alpaca + SR 5.65 6.18 / 4.43
Alpaca + NILE 5.94 6.61 / 5.10
Orca vanilla 2.20 5.61 / 4.41
Orca + SR 3.06 6.18 / 5.51
Orca + NILE 4.63 10.77 / 8.46
Qwen2.5-7B
Alpaca vanilla 7.13 13.84 / 7.83
Alpaca + SR 6.78 15.40 / 8.90
Alpaca + NILE 8.13 17.42 / 12.24
Orca vanilla 6.60 19.19 / 13.42
Orca + SR 7.05 18.45 / 14.34
Orca + NILE 7.31 20.55 / 16.58
Qwen2.5-14B
Alpaca vanilla 7.33 15.37 / 8.16
Alpaca + SR 7.73 21.56 / 12.59
Alpaca + NILE 8.06 28.55 / 17.45
Orca vanilla 7.68 19.99 / 17.44
Orca + SR 7.90 24.40 / 18.85
Orca + NILE 8.21 32.12 / 29.82
Table 5: Experiment results of more LLMs on Alpaca and OpenOrca datasets. The highest values are bolded, and the second highest is underlined. Complete results on more benchmarks are placed in Table 17.
@@ 4.8 Ablation Study
Method A.-H. $\uparrow$ A.-E. V2 $\uparrow$ MTB. $\uparrow$ BBH $\uparrow$ A.+K. w. FD 4.80 10.75 / 9.38 6.67 60.73 A.+K. w. FS 1 IKE 4.50 11.20 / 9.75 6.72 59.25 A.+K. w. FS 2 IKE 4.50 10.82 / 10.56 6.76 61.40 O.+K. w. FD 5.20 13.67 / 11.21 7.51 64.03 O.+K. w. FS 1 IKE 4.90 12.46 / 10.99 7.40 63.89 O.+K. w. FS 2 IKE 5.50 13.00 / 11.50 7.43 64.29
Table 6: Effects of IKE with different fewshot numbers (FS) in Llama-3. The highest values are bolded, and the second highest is underlined. For brevity, Alpaca + KSR and Orca + KSR are abbreviated as A.+K. and O.+K., respectively.
Method A.-H. $\uparrow$ A.-E. V2 $\uparrow$ MTB. $\uparrow$ BBH $\uparrow$ A.+N. wo. ICF 4.50 10.82 / 10.56 6.76 61.40 A.+N. w. ICF (low) 4.80 10.69 / 10.43 6.90 61.40 A.+N. w. ICF (med.) 4.30 9.99 / 9.81 6.65 61.56 A.+N. w. ICF (high) 4.50 9.92 / 9.70 6.79 61.71 O.+N. wo. ICF 5.50 13.00 / 11.50 7.43 64.29 O.+N. w. ICF (low) 6.00 13.70 / 12.11 7.48 64.05 O.+N. w. ICF (med.) 5.00 13.27 / 11.59 7.43 64.09 O.+N. w. ICF (high) 4.80 13.19 / 11.49 7.30 63.95
Table 7: Effects of ICF using different $\beta$ parameters in Llama-3. The highest values are bolded, and the second highest is underlined. Alpaca + NILE and Orca + NILE are abbreviated as A.+N. and O.+N. for simplicity.
Effects of Different Internal Knowledge Sources
@@
Method A.-H. $\uparrow$ A.-E. V2 $\uparrow$ MTB. $\uparrow$ BBH $\uparrow$ A.+K. (Mistral) 4.00 9.14 / 7.29 6.64 57.67 A.+K. (Llama) 4.80 10.75 / 9.38 6.67 60.73 O.+K. (Mistral) 5.10 12.50 / 10.25 5.93 22.32 O.+K. (Llama) 5.20 13.67 / 11.21 7.51 64.03
Table 8: Effects of KSR in Llama-3 finetuned with internal knowledge from different LLMs. The highest values are bolded. Here Alpaca + KSR and Orca + KSR are abbreviated as A.+K. and O.+K. for brevity.
Effects of IKE Fewshot Number
Table 6 examines how different few-shot numbers of demonstration learning in IKE affect LLM performance. Here we evaluate three variants: 1) w. FD, which extracts LLM’s internal knowledge with a fixed 2-shot demonstration described in A.4.2; 2) w. FS 1 IKE, which retrieves the top 1 most similar samples with BM25 as demonstrations; and 3) w. FS 2 IKE, which retrieves the top 2 most similar samples with BM25 as demonstrations; Though w. FS 2 IKE leads to degradation in some benchmarks, such as Arena-Hard for Alpaca and Alpaca-Eval for Orca, it still achieves overall improvements with BBH for Alpaca increasing by 0.7 and Arena-Hard for Orca increasing by 0.3. The results show that IKE is necessary for unaligned LLMs to more effectively extract internal knowledge, while fixed prompting reaches subpar performance.
Effects of KSR
@@
We evaluated KSR’s effectiveness in improving internal consistency between world knowledge from instructions and the model’s internal knowledge. Our experiments assessed the degree to which responses incorporated internal knowledge across various models. We compared the models’ vanilla, KSR-generated, and SR-generated responses for 10K randomly sampled instructions by calculating sentence similarity scores. As shown in Figure 3, outputs generated by KSR exhibit a similarity score distribution significantly closer to 1 compared to SR and the vanilla baseline, with Chi-squared test p-values below 0.01. Additional results in A.4.3 further validate these findings. These results strongly support the effectiveness of KSR in enhancing internal consistency by integrating world and internal knowledge.
Effects of ICF
Table 7 looks into the effect of ICF. $\beta$ is set to 1-st percentile in Alpaca + NILE w. ICF (low) and to 2-nd percentile in Orca + NILE w. ICF (low). We set $\beta$ to 5-th and 10-th percentile for NILE w. ICF (medium) and NILE w. ICF (high). The results empirically prove that striking a balance between consistent and inconsistent knowledge in the IFT dataset is necessary for NILE to achieve ideal performance. We find the general advantage of ALPACA + NILE w. ICF (low) over ALPACA + NILE w. ICF (medium) and ALPACA + NILE w. ICF (high) and discarding ICF (ALPACA + NILE wo. ICF), indicating that a surplus of overly consistent or inconsistent samples in IFT datasets both hurt LLM’s performance, and it is crucial to find the middle ground in these samples. This experiment further verifies our design choices of the ICF phase.
@@ 5 Conclusion
We present NILE, an innovative framework designed to enhance the training efficacy of IFT datasets by aligning them with LLMs’ internal knowledge. Our extensive experiments demonstrate substantial improvements across various benchmarks, highlighting the crucial role of maintaining consistency between models’ internal knowledge and external knowledge in datasets. Each component of the NILE framework has been validated, reinforcing its importance in achieving better alignment. NILE offers promising directions for boosting the capabilities of LLMs and unlocking their full potential.
@@
While NILE can already obtain satisfactory performance, future works should expand NILE’s training by utilizing the complete OpenOrca dataset rather than the current 50,000-sample subset (5% of the dataset), due to limited time and computational resources. To ensure a fair comparison of experiments in our study, we maintain a consistent dataset size by reverting to the original answer rather than discarding samples during the ICF phase. In future work, we aim to explore more advanced data selection techniques for the ICF process. Additionally, future research should examine NILE’s capability for iterative instruction refinement, as the current implementation uses only a single revision pass. These expansions could further enhance NILE’s instruction-following capabilities.
@@
We conducted this study strictly under the guidance of community ethical principles. The utilized IFT datasets are reported to be safe and free of content that may contain discrimination, personally identifiable information, or any other undesirable behaviors. We meticulously curate our instructions to the LLMs to ensure that the tasks are limited to knowledge generation and knowledge-relevant revisions, thereby avoiding content that may pose ethical concerns.
Acknowledgement
The work described in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 2410072, RGC R1015-23). As the first author, I would like to express my heartfelt gratitude to my family, co-authors, and advisor, Prof. Irwin King, for their unwavering support and invaluable guidance throughout this work.
I would also like to dedicate this acknowledgment to the memory of 方大同 (Khalil Fong), an extraordinarily talented and influential R&B singer-songwriter. During the countless days and nights spent on this research and my PhD journey, his music and life philosophy provided me with comfort and strength to overcome every obstacle. His passing on February 21, 2025, was a profound loss for Chinese pop music and for me personally. I believe his vision and influence will continue to accompany us, transcending time and mortality, into the distant future. 成長是永遠,離別是空懸。在千尋之外,我依然存在。 (Growth is everlasting, and parting is always in suspense. Beyond searches, I still exist.)
@@ A.1 Implementation Details of NILE
For all experiments in this work, we use the Python 3.10.14 environment and vLLM 0.5.5 library [^11] https://github.com/vllm-project/vllm for LLM local inference of both Mistral and Llama-3. For vLLM inference hypermeters, we set the random seed to 42, max_tokens to 1024, temperature to 0.7, top_k to 50, top_p to 0.7, and repetition_penalty to 1. We run all experiments on a server with an Intel Xeon Silver 4309Y CPU and 8 Nvidia RTX A6000 GPU having 48GB GDDR6 VRAM, and we utilize official checkpoints Mistral-7B-v0.3121212https://huggingface.co/mistralai/Mistral-7B-v0.3 for Mistral and Meta-Llama-3.1-8B [^13] https://huggingface.co/meta-llama/Llama-3.1-8B for Llama-3. For LLM instruction fine-tuning in this work, we choose llama-recipes [^14] https://github.com/meta-llama/llama-recipes for Llama-3 and alignment-handbook [^15] https://github.com/huggingface/alignment-handbook for Mistral. For Llama-3 fine-tuning, we set context_length to 2048, gradient_accumulation_step to 32, learning rate to 2e-5, and training batch size to 4. As for Mistral fine-tuning, we set context_length to 2048, gradient_accumulation_step to 32, learning rate to 2e-5, training batch size to 4, lr_scheduler_type to ”cosine”, num_train_epochs to 3, and warmup_ratio to 0.1. Fine-tuning for both Llama-3 and Mistral is done within 5 hours using 8 A6000 GPU.
@@ A.1.1 Internal Knowledge Extraction (IKE)
For demonstration sample, we randomly sample $m=5,000$ instruction pairs $q^{d}_{i}=\{\mathrm{instruction}^{d}_{i},\mathrm{input}^{d}_{i}\}$ from Alpaca dataset [^16] https://huggingface.co/datasets/vicgalle/alpaca-gpt4, since instructions in it are simple and straightforward, which is suitable for LLM demonstration learning. We leverage gpt-4-turbo-2024-04-09 through the OpenAI API for generating demonstrations given $q^{d}_{i}$ shown in Table 1. For GPT-4 endpoints, we use the OpenAI 1.42.0 Python library and set n to 1, temperature to 0.7, and max_tokens to 1,024. We stick to the regulations from the OpenAI company when accessing its API. For retriever $\mathcal{R}$ in IKE, we choose the BM25 and Contriver implementation [^17] https://github.com/castorini/pyserini from the Pyserini 0.38.0 library.